OpenStudy (anonymous):

Calculus help plz? You are given the line (L) which is the intersection of planes 3x-2y-z=-4 and x-4y-3z=2 . Find the equation of the projectionof the line L on the plane 5x+2y+2z=7

OpenStudy (anonymous):

do you have any idea? I am not sure whether I can help or not, but I think we can work together, I am studying this stuff, too

OpenStudy (anonymous):

plane P1 has normal vector n1 = (3,-2,-1) plane P2 has normal vector n2 = (1,-4,-3) the intersection line d= n1x n2 = (1,8,-10)